# MemAscend: System Memory Optimization for SSD-Offloaded LLM Fine-Tuning

### v1

@@ II-C GPU Memory Efficiency Optimizations
Although SSD offloading addresses the capacity limitations for storing large model states, optimizing the usage of limited, high-bandwidth GPU memory remains crucial for performance. Several techniques enhance GPU memory efficiency and can be used in conjunction with SSD offloading systems like ZeRO-Infinity and MemAscend. By reducing the GPU memory footprint, these methods allow the system to accommodate larger models, increase batch sizes, or enable longer context lengths, potentially improving overall training throughput.
@@ II-C1 Liger-Kernel
Liger-Kernel [8] provides a library of optimized Triton kernels for common operations in LLM training, such as RMSNorm, SwiGLU, and Cross Entropy loss computation. Standard deep learning systems such as PyTorch often execute these operations using multiple separate CUDA kernel launches, leading to intermediate memory allocations and kernel launch overhead. Liger-Kernel employs kernel fusion to combine these operations into single, optimized kernels. This fusion avoids materializing large intermediate tensors (e.g., the full logit tensor before softmax in Cross Entropy), thereby reducing peak GPU memory usage and improving execution speed. By integrating Liger-Kernel, we can maximize GPU computational efficiency and minimize memory spikes during the forward and backward passes, making better use of the GPU resources freed up by SSD offloading.
@@ II-C2 Flash-Attention
Flash-Attention [14] is an I/O-aware exact attention algorithm designed to address the quadratic memory and time complexity of standard attention mechanisms, particularly for long sequences. Instead of materializing the large $N\times N$ attention matrix in GPU high-bandwidth memory, Flash-Attention uses tiling and takes advantage of faster on-chip SRAM. Calculate the output of attention block by block, significantly reducing the number of reads and writes from memory to HBM. This approach reduces the memory complexity of attention from $O(N^{2})$ to $O(N)$ (linear in context length) without resorting to approximations, thus maintaining model precision. Utilizing Flash-Attention drastically cuts the GPU memory required for the attention mechanism, which is often a bottleneck for long-context training, and simultaneously speeds up computation. This works well with SSD offloading by minimizing the GPU memory required for dynamic activations.
@@ II-C3 Gradient Checkpointing
Gradient checkpointing [7], also known as activation recomputation or activation checkpointing, is a technique that trades additional compute during the backward pass for reduced GPU memory usage during the forward pass. Instead of storing all intermediate activations required for gradient calculation, only a subset of activations (checkpoints) is saved. During the backward pass, the activations needed between checkpoints are recomputed on-the-fly starting from the nearest preceding checkpoint. Various strategies exist, such as checkpointing every $k$ layers or selectively checkpointing based on memory usage and recomputation cost. This technique significantly lowers the maximum GPU memory required for activations, making it essential for training large models or using long sequences.
II-C4 Offload Gradient Checkpointing
Building upon gradient checkpointing, offload gradient checkpointing further reduces the footprint of the GPU memory by moving the saved checkpoint activations themselves from the GPU memory to the CPU system memory. During the backward pass, these checkpoints are transferred back to the GPU just before they are needed for recomputation. ZeRO-Infinity includes the capability to offload activation checkpoints to CPU memory. Unsloth [28] also provides an implementation focused on efficient CPU offloading using asynchronous transfers. While this technique dramatically frees up GPU memory, potentially enabling much longer context lengths or larger batch sizes, it makes the training process more sensitive to system memory capacity and bandwidth. Our work, MemAscend, by optimizing system memory usage and reducing overhead within the ZeRO-Infinity system, directly enhances the feasibility and effectiveness of offload gradient checkpointing. The system memory saved by MemAscend can be repurposed to store more or larger offloaded checkpoints, thus enabling larger trainable model sizes or context lengths than possible with the baseline ZeRO-Infinity under the same hardware constraints.
Although SSD offloading systems like ZeRO-Infinity provide a foundation for training massive models by leveraging heterogeneous memory, significant inefficiencies remain, particularly in system memory management. Our work focuses on the system memory bottlenecks inherent in the SSD offloading system itself, with integration with other related works like SSDTrain, Liger-Kernel, Flash-Attention, and offload gradient Checkpointing, and optimizing various aspects of memory. thereby improving its overall efficiency and scalability across various configurations and enabling better synergy with other optimization techniques.
@@ III Observation and Motivation
III-A System Memory Fragmentation Challenge
In an SSD offloading system, there is a system memory buffer pool fragmentation issue that occupies more system memory than necessary. This buffer is due to the fact that all model weights are stored in SSD storage and the system pre-fetches only the necessary weights during forward or backward propagation.
To facilitate this process, the system implements a buffer pool that manages all pre-allocated pinned memory. When a layer’s weights need to be fetched, the buffer pool assigns a pinned system memory buffer to it. The process involves performing I/O operations to load the weight tensor from the SSD into the pinned system memory, followed by a host-to-device copy to transfer the data to the GPU. Once the weight is no longer needed, the pinned system memory buffer is released and returned to the buffer pool. Each pinned memory buffer within the buffer pool is the same size, and the number of buffers depends on the model architecture. Since the buffer is needed every time a PyTorch submodule performs a forward operation, it is challenging to analyze statically and can dynamically change if different activation optimization algorithms are enabled (e.g., gradient checkpointing changes the frequency with which submodule within the model invokes forward operations). However, in the general case, the buffer size is designed to fit the weights of one to a few transformer blocks.
The main issue is that the buffer size is determined by the largest weight tensor, leading to internal fragmentation. Modern models typically have an embedding layer with a significantly larger weight size compared to other layers. For instance, in the Llama 3 8 billion model, the embedding layer size is determined by a vocabulary size of 128,256 multiplied by a hidden size of 5,120, whereas other layers are much smaller. For example, the second largest weights in the model are usually the up, gate, and down projection layers in the feed-forward network, with sizes of 14,336 × hidden size 5,120, which are smaller than the embedding layer. The smallest swappable layer is the key and value projection, with a size of 1,024 × hidden size 5,120. The disparity in the sizes of the tensors exacerbates fragmentation. As a result, the total buffer size required is 13.05 GiB, while the actual swappable tensor size needed at any given time is only 3.81 GiB, leading to a 70.82% fragmentation problem.
@@ III-B Inefficiency in Pinned Memory Allocation
Pinned (or page-locked) memory is essential for efficient data transfers in SSD offloading systems like ZeRO-Infinity. System memory acts as a staging ground between high-speed GPU memory and slower, high-capacity SSD storage. GPUs require data to reside in pinned memory before initiating Direct Memory Access (DMA) over PCIe. Bypassing the extra copy from pageable memory to a pinned buffer managed by the driver significantly reduces latency and CPU load. To avoid this overhead, systems often pre-allocate pinned memory buffers for frequently transferred tensors, such as model weights fetched from SSD or gradients being offloaded. Standard memory allocators in deep learning frameworks, particularly PyTorch’s CachingHostAllocator, can introduce substantial inefficiencies in this context. The allocator caches pinned memory and rounds each requested size S up to the next power of two. This alignment minimizes fragmentation in workloads with frequent, variably sized allocations and deallocations, but it becomes inefficient when applied to SSD offloading.
In SSD offloading, pinned memory allocations tend to be large and long-lived. Flat buffers for gradient partitions, staging areas for optimizer states, and large tensors such as embeddings are typically allocated once and persist throughout training. When such buffers, often hundreds of MiB or several GiB, are aligned to the next power of two, the internal fragmentation becomes significant. For instance, an allocation slightly over 2 GiB may be rounded up to 4 GiB, wasting nearly 2 GiB. The original motivation for this alignment, which is avoiding fragmentation in dynamic allocation patterns, does not apply when the allocations are large and static. Reducing this alignment overhead is critical for efficient memory use in SSD offloading systems.
(a) ZeRO-Infinity
(b) MemAscend
Figure 2: Comparison of tensor lifetimes during gradient overflow checks in SSD offloading.
@@ III-C Peak Memory Bottlenecks in System Memory
In ZeRO-Infinity, the flat buffer for gradient partitioning is a contiguous memory block that holds partitioned gradients to streamline communication and computation during distributed training. With gradient accumulation enabled, this buffer supports seamless accumulation. When ZeRO-Offload or ZeRO-Infinity is active, the buffer resides in system memory, allowing efficient transfer of temporary GPU results into it.
A key limitation, however, is that the buffer size mirrors the model parameter size. To support accumulation across iterations, the buffer typically uses fp32 even when forward and backward passes run in fp16. This becomes a major bottleneck in large-scale models, as the flat buffer size grows with model complexity. As shown in Figure 2(a), the peak memory usage primarily stems from gradient overflow checks. In mixed-precision training with fp16, overflows can occur in intermediate results, requiring a check after every iteration. If an overflow is found, the system reduces the scaling factor to prevent future instability. This check introduces two main challenges.
First, the memory overhead of detecting overflows is significant. Three conditions must be evaluated: infinity, negative infinity, and NaN. In PyTorch, this involves calling isinf() followed by any(). Internally, isinf() triggers isabs() to duplicate the tensor before performing comparisons, which creates a temporary copy and a Boolean tensor, resulting in a memory spike to about 2.25× the original size. isnan() avoids isabs() but still produces a Boolean tensor, adding another 1.25× memory peak. For example, when training an 8-billion-parameter model, the flat buffer on the CPU consumes 29.91 GiB, but the overflow check pushes usage to 67.30 GiB. Second, the sequence of operations in the overflow check introduces significant latency: (1) isabs(), (2) isinf(), (3) any(), (4) isnan(), and (5) any(). This chain of steps adds roughly 5507 ms per iteration for an 8-billion-parameter model on the Configuration 1 machine described in Table II, substantially slowing training. Since this check runs every iteration, the cumulative impact on performance is considerable.
@@ III-D Filesystem Overhead in SSD offloading
In the SSD offloading system, the SSD serves as extended memory. It stores low-precision compute weights, such as fp16, along with the full-precision fp32 master weights, gradients, and optimizer states. These data are loaded or offloaded depending on the computational context. The current implementation uses the filesystem with the O-DIRECT flag, which bypasses the page cache to ensure that data transfers go directly to the SSD. Each tensor is offloaded to a separate file, allowing the filesystem, such as ext4, to manage storage allocation without direct intervention. However, writing to ext4 carries overhead. The filesystem maintains metadata for each file, including name, size, permissions, and timestamps, which requires additional I/O and increases latency. Allocating disk blocks can also involve searching for contiguous space, updating allocation tables, and maintaining alignment, all of which add computational cost and risk fragmentation. When the SSD is used as an extension of memory and subjected to sustained I/O, these filesystem overheads accumulate. This added latency can limit the efficiency gains expected from offloading, particularly during training.
III-E Motivation
The inefficiencies described above, including buffer fragmentation, overhead from pinned memory allocation, and memory spikes during tasks like gradient overflow checks, lead to substantial system memory waste. While SSD offloading already enables training significantly larger models on machines with limited resources, as shown in Table I under a 128 GiB system memory cap, it still hits a hard ceiling. Regardless of whether the setup involves a single GPU or multiple, system memory bottlenecks inherent to SSD offloading limit scalability and undercut cost-efficiency. This work aims to eliminate those inefficiencies within the SSD offloading context. Reducing system memory waste brings clear advantages across all SSD offloading scenarios. It either allows training larger models within the same memory limits or reduces the hardware required for current models, improving cost efficiency. The saved memory space can also store more offloaded activation checkpoints, which supports longer context lengths or larger batch sizes. This not only extends the feasible model scale and input capacity, it also improves training throughput by increasing computation per I/O cycle. Addressing overheads like I/O latency and overflow check times further enhances efficiency and scalability.
@@ III-E Motivation
Type Model Peak DRAM (GiB)
@@
IV Methods
IV-A Overview of Methods
This study presents a set of optimizations to address critical memory inefficiencies in SSD offloading systems. The approach targets four primary issues: system memory fragmentation in buffer pool management and pinned memory allocation, peak memory usage due to gradient overflow checks, and file system overhead in SSD offloading. To tackle these inefficiencies, the approach employs four key optimizations. First, an adaptive buffer pool strategy reduces internal fragmentation caused by monolithic buffer allocations. Second, a custom memory allocator minimizes waste from excessive alignment. Third, a fused kernel leverages OpenMP to perform efficient gradient overflow checks, reducing memory spikes and operation latency. Fourth, a direct SSD space management scheme replaces the filesystem-based SSD offloading mechanism, eliminating metadata overhead and lowering I/O latency. The approach integrates into the ZeRO-Infinity framework, which combines SSD offloading, gradient checkpointing, and the Liger Kernel to enhance GPU memory efficiency. These optimizations improve performance, alleviate memory pressure, and enable larger batch sizes or longer context windows by freeing system memory for activation offloading.
Figure 3 illustrates the MemAscend architecture. The approach affects three distinct stages of the training process. During initialization, the system allocates various buffers, including gradient partition flat buffers, buffer pools for optimizer states, and other critical allocations. Instead of relying on PyTorch’s default pinned memory mechanism, MemAscend uses a Zero Overhead Allocation C++ kernel to reduce unnecessary alignment and fragmentation. In the layer-wise forward and backward passes, the system pre-fetches weight tensors from SSD to CPU, then transfers them to the GPU for an all-gather operation. A hook linked to this transfer activates the parameter swapper, which moves tensors from the SSD to system memory. The swapper queries the Adaptive Buffer Pool to secure a buffer matching the current layer’s size. Once allocated, the buffer manager invokes the Direct NVMe Engine to execute the required I/O operations. After loading, the data transfer to GPU memory, and the buffer is released. This cycle repeats across layers during training. In the final stage, before the optimizer step, the system verifies that accumulated gradients have not overflowed. A Fused Overflow Check Kernel performs a single-pass check with minimal memory overhead and computation, reducing redundancy.

### v2

@@ II-C GPU Memory Efficiency Optimizations
Although SSD offloading alleviates the capacity constraints of storing large model states, optimizing the usage of limited, high-bandwidth GPU memory remains essential for maintaining training performance. Several techniques have been proposed to improve GPU memory efficiency, particularly the residual memory, and can be applied alongside SSD offloading systems such as ZeRO-Infinity and MemAscend. By reducing the GPU memory footprint, these methods enable the training of larger models, support increased batch sizes, and allow for longer context lengths, potentially enhancing overall throughput. Otherwise, the SSD offloading still faces the out-of-memory (OOM) issue when increasing batch size or context lengths, since GPU memory cannot accommodate all activations and intermediate values. Detailed GPU memory usage with different memory efficiency methods enabled is shown in Figure 2. Accordingly, it can be seen that the GPU memory usage is further reduced after enabling the following optimizations alongside Zero-Infinity. Therefore, this paper integrates these methods with the state-of-the-art offloading framework (i.e., Zero-Infinity) and utilizes it as the comparison baseline for the proposed MemAscend.
(a) Context length=512
(b) Context length=32768
Figure 2: Comparison of GPU memory usage between short and long context lengths with a batch size of 4 for training an 8-billion-parameter model, across different GPU memory efficiency optimizations. The y-axis uses a base-10 logarithmic scale. GC represents enabled Gradient Checkpointing, Liger/Flash represent enabled Liger-Kernel and FlashAttention, and Offloaded-GC represents enabled Offloaded Gradient Checkpointing.
@@ II-C1 Liger-Kernel
Liger-Kernel [12] provides a library of optimized Triton kernels for common operations in LLM training, such as RMSNorm, SwiGLU, and Cross Entropy loss computation. Standard deep learning frameworks like PyTorch typically execute these operations through multiple separate CUDA kernel launches, resulting in additional intermediate memory allocations and kernel launch overhead. Liger-Kernel addresses this by employing kernel fusion, combining multiple operations into a single, optimized kernel. This eliminates the need to materialize large intermediate tensors (e.g., the full logit tensor before softmax in Cross Entropy), thereby reducing peak GPU memory usage and improving execution speed. By integrating Liger-Kernel, GPU computational efficiency is maximized, and memory spikes during forward and backward passes are minimized, making more effective use of GPU resources freed by SSD offloading.
@@ II-C2 Flash-Attention
Flash-Attention [17] is an I/O-aware, exact attention algorithm designed to overcome the quadratic time and memory complexity of standard attention mechanisms, especially for long sequences. Instead of materializing the entire $N\times N$ attention matrix in high-bandwidth GPU memory, Flash-Attention computes attention outputs block-by-block using tiling and leverages fast on-chip SRAM. This significantly reduces memory traffic to HBM and lowers memory complexity from $O(N^{2})$ to $O(N)$, all while maintaining numerical accuracy without relying on approximation. By dramatically cutting the memory required for attention and accelerating computation, Flash-Attention pairs well with SSD offloading strategies by reducing the GPU memory pressure from dynamic activations.
@@ II-C3 Gradient Checkpointing
Gradient checkpointing [10], also known as activation checkpointing or recomputation, reduces GPU memory usage during the forward pass by trading it for additional computation in the backward pass. Instead of storing all intermediate activations needed for backpropagation, only a small subset (checkpoints) is retained. During the backward pass, missing activations are recomputed on-the-fly from the nearest checkpoint. Strategies such as fixed-interval checkpointing (e.g., every $k$ layers) or adaptive checkpointing based on memory cost can be used. This technique significantly reduces peak memory consumption, enabling the training of deeper models or longer sequences.
II-C4 Offloading Checkpointed Gradients
Building on gradient checkpointing, offloading the checkpointed gradient can further reduce GPU memory consumption by transferring the checkpointed activations from GPU memory to CPU system memory. These checkpoints are transferred back to the GPU just in time for recomputation during the backward pass. ZeRO-Infinity includes support for this mechanism, and systems like Unsloth [33] provide efficient implementations using asynchronous transfers. While this technique can greatly expand feasible model sizes and context lengths, it introduces dependency on system memory bandwidth and capacity. Our work, MemAscend, enhances the practicality of offloading checkpointed gradient by reducing system memory fragmentation and overhead within ZeRO-Infinity. The memory reclaimed by MemAscend can be reallocated to store more or larger offloaded checkpoints, thereby supporting larger models or sequences under the same hardware constraints.
In summary, although SSD offloading systems like ZeRO-Infinity lay the groundwork for training massive models using heterogeneous memory, our investigation reveals that substantial inefficiencies remain, particularly in system memory management. The design goal of MemAscend is to directly address these limitations by optimizing system memory usage within SSD offloading systems. Then, through integration with complementary techniques such as SSDTrain, Liger-Kernel, Flash-Attention, and offloaded gradient checkpointing, MemAscend enhances overall memory efficiency and scalability. These improvements allow SSD offloading systems to operate effectively across a wider range of hardware configurations while enabling better synergy with other memory optimization strategies.
@@ III Observation and Motivation
This section presents key observations that reveal critical inefficiencies in existing SSD offloading systems. Despite their ability to extend memory capacity, challenges such as system memory fragmentation, inefficient pinned memory allocation, peak memory bottlenecks, and filesystem-induced I/O overhead remain unresolved. These issues degrade performance and limit scalability. The final subsection outlines the motivation for MemAscend, which aims to address these problems to enhance the efficiency of SSD offloading frameworks.
III-A System Memory Fragmentation Issue
In the SOTA SSD offloading technique, a system memory buffer pool, referred to as the parameter buffer pool, is designed to prefetch model weights from SSDs into DRAM for a brief window while each layer executes its forward or backward pass. Nonetheless, the current design of the system memory buffer pool leads to the issue of system memory fragmentation that inflates memory consumption well beyond what the models truly need. The current design equips the buffer pool with pre-allocated pinned memory buffers of the same size. Then, when a layer requires its weights, the buffer pool assigns a pinned memory buffer to it. The weights are loaded from SSD into this memory buffer via I/O operations, followed by a host-to-device copy to transfer the data to the GPU. Once the weights are no longer needed, the memory buffer is returned to the pool. Since prefetching is employed, multiple transformer blocks [^2] Transformer blocks are the fundamental building units of the transformer architecture and are designed to process input sequences by capturing relationships between elements through attention mechanisms and feed-forward networks. are typically in flight at once. The required buffer pool size, therefore, equals the size of each memory buffer multiplied by the number of memory buffers needed per transformer block, and further multiplied by the number of transformer blocks in flight.
The core issue is that the size of each pinned memory buffer in the buffer pool is fixed based on the largest weight tensor in the model, resulting in severe internal fragmentation. In modern transformer-based models, the embedding layer often contains the largest weights. For example, in LLaMA 3 8B, the embedding layer is sized based on a vocabulary of 128,256 and a hidden dimension of 5,120, making it significantly larger than other layers. In contrast, layers such as the up, gate, and down projections in the feed-forward network (typically sized 14,336 × 5,120) and the key/value projection layers (1,024 × 5,120) are much smaller. This imbalance leads to inefficient memory usage. As a result, although the total buffer pool is sized at 13.05 GiB, the actual memory required to hold the largest amount of tensors in use at any time is only 3.81 GiB, leading to 70.82% internal fragmentation.
@@ III-B Inefficiency in Pinned Memory Allocation
Pinned (page-locked) system memory underpins high-bandwidth transfers in SSD offloading frameworks. This is because GPUs can only launch PCIe DMA to physical pages that remain resident; all data shuttled between SSD and GPU DRAM must first land in pinned system memory. Pre-allocating these buffers eliminates the extra copy from pageable DRAM to a driver-managed staging area, cutting latency and CPU overhead. Consequently, large tensors that move frequently, model weights fetched from SSD, or gradients written back, are allocated once and reused for the duration of training. The default pinned-memory allocator in PyTorch, CachingHostAllocator rounds every request up to the next power of two to curb fragmentation in highly dynamic workloads. However, that policy backfires in SSD offloading, where allocations are both large and long-lived. Flat buffers for gradient shards, optimizer states, or embedding matrices may be hundreds of MiB to several GiB; aligning a 2.1 GiB request to 4 GiB needlessly wastes almost 2 GiB. Because these buffers persist, the alignment overhead turns into permanent internal fragmentation rather than a transient cost. Eliminating such over-alignment is therefore critical to reclaim system memory and raise the efficiency in SSD offloading systems.
@@ III-C Peak Memory Bottlenecks in System Memory
In ZeRO-Infinity, the flat buffer for gradient partitioning is a single contiguous block that stores all partitioned gradients, enabling streamlined communication and computation during distributed training. When gradient accumulation is enabled, the same buffer is reused across iterations. Because ZeRO-Offload and ZeRO-Infinity place the buffer in system memory, temporary GPU results can be copied into it efficiently. However, a critical drawback is that the buffer’s capacity needs to match the total model-parameter size. To support accumulation, the buffer is allocated in fp32 even when forward and backward passes use fp16, so its footprint scales directly with model complexity. As Figure 3 shows, the main peak-memory issue is the gradient-overflow check performed each iteration in mixed-precision training. If any intermediate value overflows, the scaling factor is reduced to avoid instability. This check introduces both memory and latency overhead.
Figure 3: Tensor lifetimes during gradient overflow checks in ZeRO-Infinity.
To detect overflows during the overflow check, each value must be examined to determine whether it meets any of the following three conditions: infinity, negative infinity, or NaN. In PyTorch, this involves calling isinf() followed by any(). Internally, isinf() triggers isabs() to duplicate the tensor before performing comparisons, which creates a temporary copy and a Boolean tensor, resulting in a memory spike to about 2.25× the original size, as shown in Steps 2 of Figure 3. At Step 3, isnan() avoids isabs() but still produces a Boolean tensor, adding another 1.25× memory peak. For example, when training an 8-billion-parameter model, the flat buffer on the CPU consumes 29.91 GiB, but the overflow check pushes usage to 67.30 GiB. On the other hand, the sequence of operations in the overflow check introduces significant latency: (1) isabs(), (2) isinf(), (3) any(), (4) isnan(), and (5) any(), as shown in Steps 2 to 6 in Figure 3. This chain of steps adds roughly 5507 ms per iteration for an 8-billion-parameter model on the Configuration 1 machine described in Table II, accounting for 13.36% of the single iteration latency when training with a context length of 4096 and a batch size of 8, thereby substantially slowing down the training process. Since this check runs every iteration, the cumulative impact on performance is considerable. Thus, both the memory spike and the extra computation make gradient overflow checking a major bottleneck for large-scale models.
@@ III-D Filesystem Overhead in SSD offloading
In an SSD offloading system, the SSD serves as extended memory. It stores low-precision compute weights, such as fp16, along with the full-precision fp32 master weights, gradients, and optimizer states. These data are loaded or offloaded depending on the computational context. The current implementation uses conventional file systems with the O-DIRECT flag, which bypasses the page cache to ensure that data transfers go directly to the SSD. Each tensor is offloaded to a separate file, allowing file systems, such as ext4, to manage storage allocation without direct intervention. However, writing to ext4 carries overhead. The file system maintains metadata for each file, including name, size, permissions, and timestamps, which requires additional I/O and increases latency. Allocating disk blocks can also involve searching for contiguous space, updating allocation tables, and maintaining alignment, all of which add computational cost and risk fragmentation. When the SSD is used as an extension of memory and subjected to sustained I/O, these file system overheads accumulate. This added latency can limit the efficiency gains expected from offloading, particularly during training.
@@
Type Model Peak System Memory (GiB)
@@
Figure 4: System memory overhead in the original SSD offloading system across different models. This demonstrates that the original SSD offloading system limited the trainable model size, context length, and batch size due to memory overhead.
III-E Motivation
Even though the SSD offloading technique offers a promising solution for training larger models on systems with limited resources, those observed system-memory bottlenecks intrinsic to SSD offloading lower scalability and erode cost-effectiveness. To understand the limit of existing SSD offloading techniques, a motivational experiment is conducted based on a machine with 24 GiB of GPU memory and a 128 GiB system memory limit, and results are summarized in Table I. Accordingly, the All-in-GPU approach, which stores everything on the GPU, supports training models with up to 1 billion parameters. By using ZeRO-Offload, which shifts data to CPU memory, the same system can handle models with up to 3 billion parameters. With the SOTA SSD offloading strategy, ZeRO-Infinity, it can train models with up to 8 billion parameters. It can be observed that, even though the model size is increased, the system memory usage grows more drastically and soon reaches the system limits of commodity CPUs. According to the above observations, inefficiencies such as buffer pool fragmentation, unnecessary alignment for pinned memory allocations, and memory spikes during gradient overflow checks lead to significant waste of system memory and prevent the SOTA SSD offloading strategy from accommodating larger models.
To quantify the actual required and wasted system memory usage, another motivational experiments are conducted on Configuration 2, as mentioned in Table II. As shown in Figure 4, the system wastes an average of 55.7% of its memory across different models, making system memory the next bottleneck. These limitations reduce the scalability and cost-effectiveness of SSD offloading. This study directly addresses those inefficiencies. Reclaiming wasted system memory can either increase the maximum trainable model size on current hardware or reduce the memory required for existing workloads, improving economic efficiency. The recovered capacity can also be used to store additional offloaded activation checkpoints, supporting longer context windows or larger batch sizes and boosting throughput by increasing useful computation per data movement. Finally, minimizing secondary overheads such as I/O latency and overflow check delays further enhances overall efficiency and scalability.
IV MemAscend
IV-A Overview
In this study, to resolve the critical issue of system memory inefficiencies in SSD offloading systems, MemAscend is proposed with four key optimizations, which are (1) an adaptive buffer pool, (2) an alignment-free pinned memory allocation, (3) a fused overflow check mechanism, and (4) a direct NVMe engine. Figure 5 depicts a software–hardware split, and highlighted components are implemented as part of the proposed MemAscend. The software component represents the primary offloading system written in Python, along with helper C++ kernels that handle low-level operations through a Python-C++ binding interface. This allows the offloading system to perform tasks such as memory allocation and NVMe I/O. The hardware component includes three types of memory devices used in SSD offloading systems: GPU memory, system memory, and SSD. According to the design of MemAscend, three stages of the training process are altered as follows.
During initialization, gradient-partition buffers, optimizer-state pools, and other critical regions are allocated by the alignment-free pinned memory allocation, eliminating padding waste and fragmentation of PyTorch’s default pinned memory mechanism. The forward-and-backward pass then operates in a pipeline: weights are prefetched from SSD to CPU DRAM, moved to the GPU for the all-gather operation. A hook linked to this transfer activates the parameter swapper, which moves tensors from SSD to system memory. The swapper draws exactly-sized chunks from the adaptive buffer pool and hands them to the direct NVMe engine for low-overhead I/O, releasing each chunk once its data reaches GPU memory. The above procedure cycles repeatedly across layers during training. In the final stage, before the optimizer step, a single-pass fused overflow check validates accumulated gradients with negligible extra memory or compute. Notably, the proposed MemAscend is seamlessly integrated with ZeRO-Infinity, Liger-Kernel, FlashAttention, gradient checkpointing, and offloaded gradient checkpointing, ensuring that every byte recovered by MemAscend can be redirected to larger batch sizes or longer context windows, boosting both the throughput and model capacity.
IV-B Adaptive Buffer Pool
The system memory requirements of SSD offloading frameworks are primarily driven by data transfer demands between SSDs and GPUs during layer-wise forward and backward propagation, and commonly employ asynchronous prefetching. Due to prefetching operations that will cause N number of transformer blocks to be concurrently in flight, the required buffer capacity is therefore the size of N transformer blocks. In each transformer block, since smaller tensors can remain permanently in CPU memory, the primary tensors requiring transfer include the feedforward layer weights, specifically the up, down, and gate projections, and the attention layer weights, namely the Q, K, V, and O projections. Additionally, dedicated buffers are required for the embedding layer weights and the final linear layer weights, which are often referred to as the LM head (language-modeling head) and take hidden states from the last transformer block to produce one logit for every token in the vocabulary. Thus, the overall buffering strategy needs to accommodate weights for one embedding layer, N transformer blocks each containing weight tensors, and the LM head.
To accommodate the above buffer requirements at the system memory level, a buffer pool design is included in the original SSD offloading architecture. Nonetheless, as shown in Figure 6a, since the original design fixed the buffer block size to the largest buffer component, the issue of international fragmentation emerges. In this study, as illustrated in Figure 6b, the adaptive buffer pool design is included to assign separate subpools to tensors of different shapes, avoiding the internal fragmentation that plagues the original monolithic pool in SSD offloading. According to the flow of fine-tuning, four pools are sufficient: one for embedding-related weights, one for feed-forward projections, one for the K and V projections (identical in grouped-query attention), and one for the Q and O projections. Embedding and LM-head weights share the same dimensions—vocabulary × hidden—while feed-forward projections keep the intermediate × hidden shape, so fixed pool dimensions eliminate wasted space. The resulting subgroup counts are $2$, $3N$, $2N$, and $2N$, respectively, which match the exact number of tensors and effectively remove internal fragmentation, with N denoting the number of transformer blocks that must be in flight due to prefetching. The scheme generalizes to multi-GPU deployments that employ parameter partitioning: each process loads only its shard, so per-process buffers shrink proportionally with the number of partitions, even though the overall system still reserves the same total capacity as a single-GPU setup. Thus, the adaptive buffer pool retains its efficiency across both single-card and distributed configurations.